How do you solve y=-(x-1)^2-1
Found 2 solutions by Alan3354, ankor@dixie-net.com:
Answer by Alan3354(69443)

Solve for what?

Answer by ankor@dixie-net.com(22740)

How do you solve y=-(x-1)^2-1
y = 0, FOIL (x-1)(x-1)
-(x^2 - 2x + 1) - 1 = 0
the neg changes the signs inside the brackets
-x^2 + 2x - 1 - 1 = 0
-x^2 + 2x - 2 = 0
This equation has no real root, check this using the discriminant b^2 - 4ac
2^2 - 4(-1*-2) = -8
: